OpenStudy (anonymous):

Factor completely: x2 + 6x + 8

OpenStudy (diyadiya):

(x+2)(x+4) x=-2 x=-4

OpenStudy (anonymous):

\[x^2+6x+8=x^2 + 4x+2x+8=(x+2)(x+4)\]